OpenStudy (anonymous):

solve each system by elimination Please help and solve y -2x+3y=15 -6x+6y=18

OpenStudy (anonymous):

-2x + 3y = 15 -6x + 6y = 18 Multiply equation 1 by -3 6x - 9y = -45 -6x + 6y = 18 -3y = -27 y = 9 Now substitute -2x + 3(9) = 15 -2x + 27 = 15 -2x = -12 x = 6. Hope this helps. http://www.tutorsean.net

OpenStudy (anonymous):

it does can you help me with this one too 36+7x-8y=0 -6x+6y=18

OpenStudy (anonymous):

I will come back in about an hour. I have a private tutoring session at the college in 5 minutes, but will be more than happy to help you after it. If you look at my site, there are videos on how to do different math topics, something there may help too. It's all free. http://www.tutorsean.net

OpenStudy (anonymous):

khan academy might be able to help you too, heres a elimination walkthrough, there's substitution too if you poke around a bit. good luck! :)

OpenStudy (anonymous):

36+7x-8y=0 -6x+6y=18 Okay, so first let's put the 1st equation into correct order. 7x - 8y = -36 -6x + 6y = 18 42x - 48y = -216 -48x + 48y = 144 -6x = -72 x = 12 Now plug 12 in for x into either equation: -6(12) + 6y = 18 -72 + 6y = 18 6y = 90 y = 15 So your solution is: x = 12, y = 15. Sorry it took so long to get back. I tutor on here and privately. http://www.tutorsean.net